OpenStudy (rsadhvika):

teach me indexed sets @ikram002p

OpenStudy (rsadhvika):

im not able to appreciate indexed sets on a set

OpenStudy (zzr0ck3r):

you at least need to first understand partial orders.

OpenStudy (rsadhvika):

Ohk, i can catchup i guess... i know about metric spaces and related stuff

OpenStudy (ikram002p):

indexed set is just an expression of sets that have something common related to another countable set . ok so far ?

OpenStudy (zzr0ck3r):

need not me countable

OpenStudy (zzr0ck3r):

be*

OpenStudy (ikram002p):

the indexed set need not to be countable BUT it need to be defined with respect to countable sets , i'll give examples

OpenStudy (rsadhvika):

okay so you're ordering the set based on index ?

OpenStudy (zzr0ck3r):

as the most trivial example: \(\cup_{x\in \mathbb{R}}x=\mathbb{R}\)

OpenStudy (rsadhvika):

x is the index here ?

OpenStudy (ikram002p):

this is not indexed :O , this is like union of countable infinite indexed sets

OpenStudy (rsadhvika):

i thought you could only use natural numbers as index

OpenStudy (ikram002p):

u can use rational , natural or what ever it is hmm

OpenStudy (zzr0ck3r):

any set with a total partial order can be an indexed set

OpenStudy (rsadhvika):

what is partial order ?

OpenStudy (zzr0ck3r):

http://en.wikipedia.org/wiki/Index_set 3rd example

OpenStudy (zzr0ck3r):

its basically what you think of as \(\le\)

OpenStudy (ikram002p):

ok @zzr0ck3r teach him :)

OpenStudy (zzr0ck3r):

google it.... read the one paragraph on wiki.

OpenStudy (rsadhvika):

that looks like a bijection ?

OpenStudy (rsadhvika):

looks you're pointing elements of one set from another set, based on that definition the index need not be a natural number i see

OpenStudy (zzr0ck3r):

what does?

OpenStudy (zzr0ck3r):

Its not function

OpenStudy (rsadhvika):

tje link u gave has this an index set is a set whose members label (or index) members of another set.[1] For instance, if the elements of a set A may be indexed or labeled by means of a set J, then J is an index set. The indexing consists of a surjective function from J onto A and the indexed collection is typically called an (indexed) family,

OpenStudy (zzr0ck3r):

a partially ordered set, is a set A with a relation < where three properties need to occur 1) a<a (reflexive) 2)a<b and b<a implies a=b (anti symmetric) 3) a<b and b< c implies a<c (transitive) one example would be the naturals with "divides"

OpenStudy (zzr0ck3r):

ahh I thought you were talking about < a surjection is an onto map, a bijection is a one to one and onto map

OpenStudy (rsadhvika):

got it, and also got the definition of partially ordered set

OpenStudy (zzr0ck3r):

think of a sequence, we use N to be the index. but a subsequence of some sequence would have a subset of N as the domain, so we need not have our map be onto

OpenStudy (rsadhvika):

Oh so a sequence definition also uses indexed sets is it ?

OpenStudy (zzr0ck3r):

but the indexed set need not be countable. This shows up more in set theory/topology.

OpenStudy (rsadhvika):

\[\{a_n\}\] is \(n\) an index set of natural numbers ?

OpenStudy (zzr0ck3r):

well its using an indexed set N to "index" the elements of the sequence

OpenStudy (rsadhvika):

thanks i have an example if u have time.. the one you, me, chris and eliassaab have worked sometime back... one sec

OpenStudy (zzr0ck3r):

the sequence itself could be thought of as an indexed set.

OpenStudy (zzr0ck3r):

eliassaab is very strong in this topic. I would listen to what he said. I was not trying to cut you off @ikram002p I just saw someone telling this same user the same thing yesterday, so I wanted to make sure it was clear. I was just knocked off points for assuming my index set I was countable.

OpenStudy (zzr0ck3r):

like 3 days ago, so its fresh:P

OpenStudy (ikram002p):

:D well i rather to learn instead of teaching :P i made a mistake i guess by saying countable pardon me .

OpenStudy (zzr0ck3r):

same here. On a damn test.

OpenStudy (rsadhvika):

what does it mean by taking "union of an infinite indexed family in T" ? \[\rm A \subset X , T=2^A \cup \{X\}\]

OpenStudy (ikram002p):

it means if all sets ( that follows some rule of indexed ) are in T then show there union is also in T

OpenStudy (zzr0ck3r):

family is just another word for sets, and it usually indicates that the elements of the set are themselves sets. They may use the word collection in the same manner. Infinite means there are infinite of these elements (sets) in the family(set), and they are indexed by some set (T not sure without more). So if you pick an element in T, then it points to a set inside this infinite family.

OpenStudy (zzr0ck3r):

Are they saying that T is the indexed set, or that T is a collection, and there is a subset of T of which we are to take the union of.

OpenStudy (ikram002p):

what do u mean by ? . So if you pick an element in T, then it points to a set inside this infinite family.

OpenStudy (zzr0ck3r):

If T is the indexed set

OpenStudy (zzr0ck3r):

But I think I was reading it wrong

OpenStudy (ikram002p):

nope its not

OpenStudy (zzr0ck3r):

I would need more context, but I would bet @ikram002p is right.

OpenStudy (rsadhvika):

sorry let me rephrase it again : i have a set defined as below \[\rm T=2^A \cup \{X\}\] where \(A \subset X \) and \(X\ne \varnothing\)

OpenStudy (ikram002p):

the point is , if there is a family of indexed set that all of its elements belong to T then the union is also in T

OpenStudy (rsadhvika):

and the question goes like this `Show that union of an infinite indexed family in T is also in T.`

OpenStudy (ikram002p):

yep , its the third condition to define a topology

OpenStudy (zzr0ck3r):

oh yeah what ^^ said

OpenStudy (rsadhvika):

so should i first pick some set for indexing elements in T ?

OpenStudy (zzr0ck3r):

so 2^A is the continuum cardinality after A?

OpenStudy (rsadhvika):

2^A is powerset of A

OpenStudy (rsadhvika):

\(\mathcal{P(A)}\)

OpenStudy (zzr0ck3r):

that is not how we denote the power set of A

OpenStudy (zzr0ck3r):

but it is how we denote the cardinality of the power set of A

OpenStudy (zzr0ck3r):

well it only makes sense if its a set, so I guess I just have never seen this notation used to mean power set

OpenStudy (rsadhvika):

ohk.. \[\mathcal{ T=P(A) \cup \{X\}}\]

OpenStudy (ikram002p):

dot pick up , just use the definition , \(\left \{ V_\alpha : \forall \alpha \in \Delta \right \}\in T \) and using P(A) for power set is much better

OpenStudy (ikram002p):

so by this i picked any indexed set , which all of its elements belong to T , ok so far ?

OpenStudy (rsadhvika):

I see \(\Delta \) is a set, what does it contain ?

OpenStudy (zzr0ck3r):

I am not sure here, it screams Zorns lemma to me so I may be blinded. Ill watch and learn:)

OpenStudy (rsadhvika):

im only looking for understanding indexed sets from analysis point of view i don't think im ready for topology yet :O

OpenStudy (zzr0ck3r):

This question actually seems quite trivial. Take an infinite union of elements in T, these elements are either subsets of A or they are all of X. If we take the union we will have two cases 1) X was one of our elements, then the union is X, and X is for sure an element in T 2) X was not one of our elements, then the union is a subset of A, so for sure its in T

OpenStudy (zzr0ck3r):

(because the union of subsets of A, is a subset of A)

OpenStudy (rsadhvika):

\[\large \mathcal {\bigcup_{i=1}^{\infty}U_i}\] ?

OpenStudy (zzr0ck3r):

right, then either U_i = X for some i, or not. if it does....

OpenStudy (rsadhvika):

Show that union of an infinite `"indexed family in T"` is also in T. i don't get that phrase yet but ur solution makes sense!

OpenStudy (zzr0ck3r):

well, I would say \(i\in I\) \(U_{i\in I}u_i\) saying it like you did is assuming its countable, because we use that notation to mean indexed by the reals

OpenStudy (zzr0ck3r):

err indexed by the naturals i mean

OpenStudy (rsadhvika):

can i read it as Show that `"infinite union of elements in T"` is also in T.

OpenStudy (zzr0ck3r):

yes

OpenStudy (ikram002p):

-.- its not topology view lol ok how u define a set that there elements are sets , and these sets (X_n) are intervals such that X_n =(0,1/n ) and n \in N just make a neat notation thats all

OpenStudy (zzr0ck3r):

who said anything about intervals?

OpenStudy (zzr0ck3r):

yes you can read it like that @rsadhvika T is a "family" or "collection" because its elements are themselves sets

OpenStudy (ikram002p):

its just an example xD sorry if i mixed things up

OpenStudy (ikram002p):

idk what were talking about anymore , explaining the indexed or solving the question

OpenStudy (zzr0ck3r):

ok so what I am I missing that the following does not work Let \(\large \cup_{i\in I}x_i\) be a family of indexed sets, where \(I \) is an infinite indexed set. case 1) \(X=x_i\) for some \(i\in I\) then \(\large \cup_{i\in I}x_i=X\in T\) case 2) \(X\ne x_i \forall i\in I\) then \(\large \cup_{i\in I}x_i\subseteq A\) so \(\large \cup_{i\in I}x_i\in T\)

OpenStudy (ikram002p):

xD why T involved

OpenStudy (ikram002p):

ok lets forget the T example for a second , @rsadhvika u got what indexed family means ?

OpenStudy (rsadhvika):

are u refering to \[\large \cup_{i\in I}x_i\] ?

OpenStudy (zzr0ck3r):

Why T involved????? The question defines T.

OpenStudy (ikram002p):

ok im completly ignored nvm guys

OpenStudy (zzr0ck3r):

lol I was just trying to solve the question that he posted from the other day. Not but in any more about indexed sets.

OpenStudy (rsadhvika):

i think i get it, you're indexing the elements, \(x\)'s, using the elements from infinite set \(I\)

OpenStudy (zzr0ck3r):

I will not but in any more*

OpenStudy (zzr0ck3r):

yes, and remember the x's are sets.

OpenStudy (zzr0ck3r):

they are either subsets of A or all of X

OpenStudy (rsadhvika):

okay so there is a surjective mapping from I to T

OpenStudy (rsadhvika):

why arent you using natural numbers as index set ? natural numbers are also infinite so they can index infinite sets right ?

OpenStudy (rsadhvika):

\(\large \cup_{i\in I}x_i\) is a family of indexed sets because each \(x\) is a subset. got it thanks :)

OpenStudy (rsadhvika):

\(\large \cup_{i\in I}x_i\) is a family of oranges if each \(x\) is an orange

OpenStudy (rsadhvika):

explicit definition of indexing set seems unnecessary but i get the general idea thanks to both of you!

OpenStudy (zzr0ck3r):

here is an easy example where our indexed set is \(\mathbb{R}\) Consider \[D:=\{\{x\} | \ x\in \mathbb{R}\}\] This is a set of singletons (some call it a family or collection), for each \(x\in \mathbb{R}\) we have a set \(\{x\}\) \(D\) is indexed by \(\mathbb{R}\).